OpenStudy (anonymous):

help please! how do I solve this equation by completing the square? 2x^2 - 2x -7 = 0 I can do it if there is no coefficient for the x^2 :(

OpenStudy (anonymous):

you need to divide by 2

OpenStudy (anonymous):

\[2x^2-2x=7\] \[x^2-x=\frac{7}{2}\] \[(x-1)^2=\frac{7}{2}+1=\frac{9}{2}\] \[x-1=\pm\sqrt{\frac{9}{2}}=\pm\frac{3}{\sqrt{2}}\]

OpenStudy (anonymous):

finish by adding 1 to both sides to solve for \(x\)

OpenStudy (anonymous):

oh okay so this technique is still completing the square?

OpenStudy (anonymous):

thanks... will study this thank you :D